For case 1) above (single coil), write an expression relating the slope of a B vs I plot in terms to μ0, N, and R. You will use this expression to calculate the value of the constant μ0 later on.
Answer #1

Since you have not given the name of specific coil or any diagram I am assuming it to be the toroid coil of mean radius R.

For a toroid coil the magnetic field at a mean distance R is given as

\oint Bdl = \mu_0 N I

B\oint dl = \mu_0 N I

B(2\pi R)= \mu_0 N I

B= \frac{\mu_0 NI}{2\pi R}

So the slope of the curve between B vs I

slope = \frac{\mu_0 N}{2\pi R}